# Large Language Models for Anomaly and Out-of-Distribution Detection: A Survey

### v2

@@ Appendix D Out-of-Distribution Detection Research Roadmap
D.1 Traditional Methods
Classification-based Methods.
Classification-based OOD detection methods rely on the outputs of neural networks , typically using the softmax probabilities of a classifier to determine whether a sample is in-distribution (ID) or OOD. The most common baseline is the Maximum Softmax Probability (MSP) method, which flags samples with lower softmax scores as OOD Hendrycks and Gimpel (2017). This has led to more advanced techniques that either post-process the classification outputs or modify the training process to improve OOD detection performance Liu et al. (2020). Given its alignment with classification tasks, this approach remains one of the most prominent methods for OOD detection.
Density-based Methods.
Density-based methods explicitly model the distribution of in-distribution (ID) data using probabilistic models, assuming that OOD samples will lie in low-density regions Zong et al. (2018); Abati et al. (2019); Pidhorskyi et al. (2018); Deecke et al. (2018); Sabokrou et al. (2018). Techniques such as class-conditional Gaussian models allow for probabilistic modeling of ID classes, while flow-based models are also used for density estimation Kobyzev et al. (2020); Zisselman and Tamar (2020); Kingma and Dhariwal (2018); Van Oord et al. (2016). However, density-based methods sometimes assign higher likelihoods to OOD samples, leading to challenges in reliability. Methods such as likelihood ratio-based approaches and ensembles have been proposed to address these issues, though these approaches tend to be computationally expensive and often fall behind classification-based methods in performance.
Distance-based Methods.
Distance-based methods operate under the assumption that OOD samples are farther from the centroids or prototypes of in-distribution classes in feature space. A popular parametric approach is to use Mahalanobis distance to compute the distance between test samples and class centroids Lee et al. (2018), whereas non-parametric methods are increasingly favored for their flexibility and simplicity Sun et al. (2022). These methods use various distance metrics—such as Euclidean distance, cosine similarity, and geodesic distance—to detect OOD samples.
Reconstruction-based Methods.
Reconstruc-
@@ D.2 LLM-based OOD Detection
D.3 Comparison of Traditional vs. LLM-based OOD Detection
@@
In conclusion, the transition from traditional OOD detection methods to LLM-based approaches represents a shift from task-specific training and rigid models to more flexible, generalizable systems that can handle a wider variety of tasks with minimal additional training. As research continues, a hybrid approach combining the efficiency of traditional methods with the flexibility and generalization of LLM-based models may offer the most robust solution for diverse OOD detection challenges.
Appendix E Datasets
E.1 Anomaly Detection
The field of anomaly detection using Large Language Models (LLMs) is still in its infancy, and as a result, no standardized benchmarks have been established. Nonetheless, several important datasets have emerged for specific applications. In industrial settings, datasets such as MVTec-AD Bergmann et al. (2019) and VisA  Zou et al. (2022) are widely used for image anomaly detection and localization tasks. Medical image datasets, such as Chest X-ray  Kermany et al. (2018), Head CT Felipe (2018), have been applied for detecting and localizing anomalies in medical imaging. Additionally, various other datasets are employed across a range of anomaly detection applications, including logical anomaly detection using MVTec LOCO Bergmann et al. (2022), video anomaly detection with the UCF-Crime dataset Sultani et al. (2018). For time series and tabular anomaly detection, datasets like the ODDS dataset  Rayana (2016) are common, alongside frequently used datasets such as YAHOO, ECG, SVDB, and IOPS, which focus on anomalies in monitoring systems and ECG recordings Paparrizos et al. (2022).
E.2 OOD Detection
Standard benchmarks use ImageNet  Deng et al. (2009) as the in-distribution (ID) dataset and other datasets as OOD samples Van Horn et al. (2018); Zhou et al. (2017); Xiao et al. (2010); Cimpoi et al. (2014), but the semantic gap between ID and OOD data can make detection easier. To address this, recent approaches split ImageNet classes into ID and OOD categories, with datasets like ImageNet-20 Ming et al. (2022) and ImageNet-10 Ming et al. (2022).
Beyond images, text-based OOD detection is also growing. Common datasets include 20 Newsgroups (20NG)  Lang (1995) for topic classification and SST-2  Socher et al. (2013) for sentiment analysis. Typically, 20NG and SST-2 serve as ID datasets, while the others are used as OOD. In intent detection, the CLINC150 dataset Repository (2020) is commonly used, splitting domains like Banking and Travel into ID and OOD categories.
Appendix F Quantitative Analysis and Comparison
@@ F.2 OOD Detection
Appendix G Evaluation Metrics
@@
AUROC (Area Under the Receiver Operating Characteristic curve):
This metric evaluates a detector’s overall ability to distinguish between ID or normal and OOD or anomalous samples. The ROC curve plots the true positive rate (TPR) against the false positive rate (FPR), where:
@@
AUPR (Area Under the Precision-Recall curve):
The AUPR metric is particularly useful for cases where there is class imbalance, as AUROC can be biased in such situations. The Precision-Recall curve plots precision against recall, where:
@@
FPR@N (False Positive Rate at TPR = N%):
This metric evaluates the probability of misclassifying an OOD or anomalous sample as ID or normal when the true positive rate (TPR) is set at a specified value, commonly 90% or 95%. This is crucial for real-world deployments where achieving high accuracy on ID samples is important, while also minimizing false positives for OOD or anomaly detection.
F1 Score:
The F1 score is a harmonic mean of precision and recall, providing a balanced evaluation of a model’s performance across both metrics. It is particularly useful in scenarios where there is an imbalance between the positive and negative classes, as it gives a single metric that reflects both false positives and false negatives.
@@
Appendix H General Guidelines
@@
The choice of approach is strongly influenced by the type of data being analyzed. For textual data, prompting-based methods may not always offer meaningful interpretations of anomalies or OOD detection, particularly when trying to understand patterns in semantic spaces. In these cases, generating embeddings from large language models (LLMs) and applying specialized post-hoc detection techniques can lead to better results. Fine-tuning LLMs to produce more relevant embeddings may further enhance detection accuracy. In the case of numerical data, such as time series or tabular data, prompting-based methods have been explored, though they often require carefully designed prompts or fine-tuning to capture the underlying structure of the data. For vision data, including images and videos, the development of multimodal LLMs offers greater flexibility. Both prompting-based and contrasting learning-based methods can be highly effective, as they are capable of handling the diverse characteristics of multimodal data.
@@
In fields where trust, transparency, and interpretability are critical, explanation plays a pivotal role in model selection. LLMs offer a unique advantage by not only detecting anomalies but also generating human-like explanations that provide deeper insights into the nature of the detected anomalies. This capability is especially valuable in domains where actionable insights and interpretable results are essential for decision-making. Research should focus on enhancing the ability of LLMs to explain their outputs in a way that aligns with domain-specific requirements for clarity and transparency.
@@
The optimization strategy is highly dependent on the specific data modality and the nature of the detection task. For prompting-based methods, parameter-efficient tuning techniques—such as Low-Rank Adaptation (LoRA) or adapter tuning—are essential for improving model performance without incurring high computational costs. These techniques enable models to adapt to new tasks efficiently while maintaining their generalization capabilities. Additionally, a hybrid approach where LLMs are used as generation tools alongside specialized anomaly or OOD detection algorithms can strike a balance between performance and efficiency. This allows the broad generalization capabilities of LLMs to be leveraged while optimizing for specific domain-related tasks, leading to improved accuracy and reduced computational overhead.

### v3

@@ Appendix C Out-of-Distribution Detection Research Roadmap
C.1 Traditional Methods
Classification-based Methods. Classification-based OOD detection methods rely on the outputs of neural networks , typically using the softmax probabilities of a classifier to determine whether a sample is in-distribution (ID) or OOD. The most common baseline is the Maximum Softmax Probability (MSP) method, which flags samples with lower softmax scores as OOD Hendrycks and Gimpel (2017). This has led to more advanced techniques that either post-process the classification outputs or modify the training process to improve OOD detection performance Liu et al. (2020). Given its alignment with classification tasks, this approach remains one of the most prominent methods for OOD detection.
Density-based Methods. Density-based methods explicitly model the distribution of in-distribution (ID) data using probabilistic models, assuming that OOD samples will lie in low-density regions Zong et al. (2018); Abati et al. (2019); Pidhorskyi et al. (2018); Deecke et al. (2018); Sabokrou et al. (2018). Techniques such as class-conditional Gaussian models allow for probabilistic modeling of ID classes, while flow-based models are also used for density estimation Kobyzev et al. (2020); Zisselman and Tamar (2020); Kingma and Dhariwal (2018); Van Oord et al. (2016). However, density-based methods sometimes assign higher likelihoods to OOD samples, leading to challenges in reliability. Methods such as likelihood ratio-based approaches and ensembles have been proposed to address these issues, though these approaches tend to be computationally expensive and often fall behind classification-based methods in performance.
Distance-based Methods. Distance-based methods operate under the assumption that OOD samples are farther from the centroids or prototypes of in-distribution classes in feature space. A popular parametric approach is to use Mahalanobis distance to compute the distance between test samples and class centroids Lee et al. (2018), whereas non-parametric methods are increasingly favored for their flexibility and simplicity Sun et al. (2022). These methods use various distance metrics—such as Euclidean distance and geodesic distance—to detect OOD samples.
Reconstruction-based Methods. Reconstruc-
@@ C.2 LLM-based OOD Detection
C.3 Comparison of Traditional vs. LLM-based OOD Detection
@@
In conclusion, the transition from traditional OOD detection methods to LLM-based approaches represents a shift from task-specific training and rigid models to more flexible, generalizable systems that can handle a wider variety of tasks with minimal additional training. As research continues, a hybrid approach combining the efficiency of traditional methods with the flexibility LLM-based models may offer the most robust solution for diverse OOD detection challenges.
Appendix D Quantitative Analysis and Comparison
@@
D.1 Anomaly Detection
D.2 OOD Detection
Appendix E Evaluation Metrics
@@
AUROC (Area Under the Receiver Operating Characteristic curve): This metric evaluates a detector’s overall ability to distinguish between ID or normal and OOD or anomalous samples. The ROC curve plots the true positive rate (TPR) against the false positive rate (FPR), where:
@@
AUPR (Area Under the Precision-Recall curve): The AUPR metric is particularly useful for cases where there is class imbalance, as AUROC can be biased in such situations. The Precision-Recall curve plots precision against recall, where:
@@
FPR@N (False Positive Rate at TPR = N%): This metric evaluates the probability of misclassifying an OOD or anomalous sample as ID or normal when the true positive rate (TPR) is set at a specified value, commonly 90% or 95%. This is crucial for real-world deployments where achieving high accuracy on ID samples is important, while also minimizing false positives for OOD or anomaly detection.
F1 Score: The F1 score is a harmonic mean of precision and recall, providing a balanced evaluation of a model’s performance across both metrics. It is particularly useful in scenarios where there is an imbalance between the positive and negative classes, as it gives a single metric that reflects both false positives and false negatives.
@@
Appendix F General Guidelines
@@
The choice of approach is strongly influenced by the type of data being analyzed. For textual data, prompting-based methods may not always offer meaningful interpretations of anomalies or OOD detection, particularly when trying to understand patterns in semantic spaces. In these cases, generating embeddings from LLMs and applying specialized post-hoc detection techniques can lead to better results. Fine-tuning LLMs to produce more relevant embeddings may further enhance detection accuracy. In the case of numerical data, such as time series or tabular data, prompting-based methods have been explored, though they often require carefully designed prompts or fine-tuning to capture the underlying structure of the data. For vision data, including images and videos, the development of multimodal LLMs offers greater flexibility. Both prompting-based and contrasting-based methods can be highly effective, as they are capable of handling the diverse characteristics of multimodal data.
@@
In fields where transparency and interpretability are critical, explanation plays a pivotal role in model selection. LLMs offer a unique advantage by not only detecting anomalies but also generating human-like explanations. This capability is especially valuable in domains where actionable insights and interpretable results are essential for decision-making. Research should focus on enhancing the ability of LLMs to explain their outputs in a way that aligns with domain-specific requirements for clarity and transparency.
@@
The optimization strategy is highly dependent on the specific data modality and the nature of the detection task. For prompting-based methods, parameter-efficient tuning techniques—such as Low-Rank Adaptation are essential for improving model performance without incurring high computational costs. These techniques enable models to adapt to new tasks efficiently while maintaining their generalization capabilities. This allows the broad generalization capabilities of LLMs to be leveraged while optimizing for specific domain-related tasks, leading to improved accuracy and reduced computational overhead.